What is the slope of (0,1) & (-2,0)
Answer by stanbon(75887)

What is the slope of (0,1) & (-2,0)
The slope of the line thru those two points is:
m = change in y/change in x = (1-0)/(0--2) = 1/2
============================================
cheers,
stan H.